A solution of hydrogen peroxide, H2O2, is 30.0% by mass and has a density of 11110g/L solution. Calculate the (a) molality, (b) molarity, and (c) mole fraction.


1
Expert's answer
2022-02-25T11:57:01-0500

Assume that the mass of H2O2 solution is 100 g.

Thus:

Mass of H2O= (%H2O2 × Mass of solution) / 100% = (30.0% × 100 g) / 100% = 30.0 g

Moles of H2O= Mass of H2O2 / Molar mass of H2O2 = 30.0 g / 34.0147 g mol-1 = 0.88197 mol

Solution volume = Mass of solution / Density of solution = 100 g / 1.135 g cm-3 = 88.1057 cm-3

Solution volume = 88.1057 cm-3 = 88.1057 mL = 0.088106 L


Molarity of H2O= Moles of H2O2 / Solution volume

Molarity of H2O= 0.88197 mol / 0.088106 L = 10.01 mol/L = 10.0 M

Molarity of H2O= 10.0 M


OR (short form solution):

(1.135 g soln/1 mL soln) × (30 g H2O2/100 g soln) × (1 mol H2O2/34 g H2O2) × (1000 mL/1 L) = 10.0 M


Answer: The molarity of a 30.0% hydrogen peroxide solution is 10.0 M.